given g(t)=(t+5}/{t-1}; evaluate a. g(5) and b. g(2s-5)
my ans is:
a. 5/2
b.-5
is this correct?

Answer by Alan3354(69443)

given g(t)=(t+5}/{t-1}; evaluate a. g(5) and b. g(2s-5)
my ans is:
a. 5/2
That's correct.
-----
b.-5
--
g(2s-5) = (2s-5+5)/(2s-5-1)
= 2s/(2s-6)
= s/(s-3)
----
How did you get -5 ?